OpenStudy (anonymous):

The mother is 40 years old. Her oldest daughter is 12 years old. When the mother will be twice as old as her daughter?

OpenStudy (anonymous):

(40+x) = 2(12+x) x = 16 in 16 years mom: 40 + 16 = 56 daughter: 12+16 = 28

OpenStudy (anonymous):

40+x=2(x+12) 40+x=2x+24 x=16 so when mother is 56 daughter will be 28